Tohban Report 2014-02-26

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 19 Feb 2014
End Date: 26 Feb 2014
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Nicole Duncan
List all reports


Solar Activity

The biggest news of the week was an X4.9 flare that occurred on Feb. 25 at ~00:39UT. This flare arose from AR 11990, in the SE region of the disk. This AR has not produced much since then, and in fact the Sun mostly quieted down after this, with GOES baseline around ~C1 level. There are now 9 ARs on the disk. Moderate/quiet activity is expected to continue, but there is a possibility of a big flare.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    33     4     1

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    21     4     1

There were RHESSI flares/GOES flares 302 / 38 over the time range 19-Feb-14 26-Feb-14

Memory Management

We did not make any changes to onboard settings in response to the big flare. We're still in Normal/Vigorous decimation, and taking rear events at night but not fronts. In the pass set immediately following the flare we reached a minimum SSR of ~5% and a max prior to the set of ~27%. We are now emptying at the end of the pass set, so there are no concerns about memory.

Spacecraft Status

The cold plate 1 and 2 are at 117-119K, putting them in the yellow zone.

Data Gaps


Detector issues

Detector 2 segmentation is proceeding nicely, with rear resets and front rates holding up consistently at all times except during the SAA. High-energy events were noticeable in the rear segment for the X flare.

Detectors 8 and 9 show extremely high fast rates, with the rise starting around Feb. 17th and the problem being quite bad the last several days. (Data in that interval will not be good from at least D9, and maybe D8.) David suggests that a noise problem in D9 is being picked up by D8, causing the mimicking behavior. To test whether we can fix both problems by adjusting D9, the HV on D9 will be dropped for one orbit to assess the situation, then raised on the next pass. This "experiment" will take place in the wee hours of the night tonight...

Other notes

SPD abstract deadline is Monday!!!!!

Spacecraft Management

No operations this week

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Disabled (4 minutes before and after eclipse)
Night time data (rears) Enabled
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? D8 and D9 very high fast rates
Personal tools